OpenStudy (anonymous):

James has a job with a current salary of $42,000. He wants to earn a college degree that will take him 2 years to complete and it will cost $28,000 per year. He will go to school full time and will not be able to work. Upon graduation, James' new salary will be $56,000 per year. How many years will it take James to recover his investment?

OpenStudy (tkhunny):

2*28000 = ?? (56000-42000)*(What) = ??

OpenStudy (anonymous):

what

OpenStudy (tkhunny):

Answer the question marks. They are the same.

OpenStudy (anonymous):

how many years

OpenStudy (tkhunny):

How many hears is the "What". You must answer the question marks.

OpenStudy (anonymous):

its 56000

OpenStudy (tkhunny):

Actually, I just saw the full-time and unable to work. Cost of college: 2*28000 Cost of not working: 2*42000 Total Cost of the adventure: 56000 + 84000 = 140000 Differential with better job? 56000 - 42000 = 14000 Looks like 10 years, no?